OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y=secx (pi/3,2)

OpenStudy (anonymous):

y' = secx tanx the reference angle of pi/3 is 60 degrees. cos(60) = 1/2, thus, sec(60) = 2 tan(60) = √3, so y' = 2√(3) (pi/3, 2) y = mx + b 2 = 2√(3) (pi/3) + b b = 2 - (2pi√(3))/3 y = 2√(3)x + 2 - (2pi√(3))/3 <== answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np my sleek-feathered one